OpenStudy (anonymous):

How to Simplify 5 v112

OpenStudy (anonymous):

\[5\sqrt{112}

OpenStudy (anonymous):

is that a division problem? 5 divide square root of 112?

OpenStudy (anonymous):

\[\sqrt[5{112}\]

Parth (parthkohli):

\[\Large \sqrt{112} = 4\sqrt{7} \] \[\Large (5*4)\sqrt{7} \] \[\Large 20\sqrt{7} \]

OpenStudy (anonymous):

\[5\sqrt[3{162}\]